OpenStudy (anonymous):

Write an equation for the line containing the points (0,6) and (-2,10)

OpenStudy (anonymous):

find slope using m=y2-y1/x2-1x then use slope and point (x,y) to plug into y=mx+b and solve for b then lastly plug m and b back in

OpenStudy (anonymous):

10-6/-2-0=4/-2 =-2

OpenStudy (anonymous):

using (0,6) for y=mx+ b 6=-2(0) + b 6=0 + b so b=6

OpenStudy (anonymous):

y=-2x+6

OpenStudy (anonymous):

u can check by plugging both points into to equation to see if it is true

OpenStudy (anonymous):

okay thanks again much apreciated

OpenStudy (anonymous):

welcome...anything else?

OpenStudy (anonymous):

actually that about it but thanks

OpenStudy (anonymous):

ok